Use Excel & show work - The Dynaco Manufacturing Company produces a particular product in an assembly line operation. One of the machines on the line is a drill press that has a single assembly line feeding into it. A partially completed unit arrives at the press to be worked on every 7.5 minutes, on average. The machine operator can process an average of 10 parts per hour. Determine the average number of parts waiting to be worked on, the percentage of time the operator is working and the percentage of time the machine is idle.
Answer #1

The manufacturing system represents single server queuing model.

Excel Model for the problem is as follows:

Dynaco Manufacturing Company - Single Server Queuing Model

Inter-arrival time

7.5

minutes per unit

Arrival rate

λ

8

units per hour

Serivce rate

µ

10

units per hour

Utilization of server

ρ

0.8

Average number of requests that will be waiting for service

Lq

3.20

Units

average waiting time in hour before service begins

Wq

0.40

Hours per unit

average waiting time in minutes before service begins

Wq

24.00

Minutes

Percentage of time the operator is busy

80.00%

Percentage of time the operator or machine is idle

20.00%

Excel Formula:

Dynaco Manufacturing Company - Single Server Queuing Model

Inter-arrival time

7.5

minutes per unit

Arrival rate

λ

=60/C2

units per hour

Serivce rate

µ

10

units per hour

Utilization of server

ρ

=+λ/µ

Average number of requests that will be waiting for service

Lq

=λ^2/(µ*(µ-λ))

Units

average waiting time in hour before service begins

Wq

=+C6/λ

Hours per unit

average waiting time in minutes before service begins

Wq

=+C7*60

Minutes

Percentage of time the operator is busy

=+ρ

Percentage of time the operator or machine is idle

=1-ρ

Snapshot:

1 Dynaco Manufacturing Company - Single Server Queuing Model 2 Inter-arrival time 3 Arrival rate 4 Serivce rate 5 Utilization

the average number of parts waiting to be worked on: 24 minutes

the percentage of time the operator is working: 80%

and the percentage of time the machine is idle.: 20%